A flywheel is rotating at 30rev/s. What is the total angle in radians, through which a point on the flywheel rotates in 40 s? show the Given, Unknown, Equation and Solution

1
Expert's answer
2022-01-10T09:10:13-0500

Given: "\omega=30\ \dfrac{rev}{s}" is angular velocity of the flywheel, "t=40\ s" is the time.

Unknown: "\theta", the total angle in radians, through which a point on the flywheel rotates in "t=40\ s".

Equation:


"\omega=\dfrac{\theta}{t}."

Solution:


"\theta=\omega t,""\theta=30\ \dfrac{rev}{s}\times\dfrac{2\pi\ rad}{1\ rev}\times40\ s=7540\ rad."